### v1

@@ 1 Introduction
Multibody dynamics simulation is relevant in engineering design and scientific computation, with applications in numerous fields, e.g., robotics featherstone2008rigid , vehicle and train dynamics shabana2013dynamics , biomechanics delp2007opensim , and molecular dynamics schlick2010molecular , to name a few. Traditionally, modeling multibody systems has relied on first-principles approaches such as Newton-Euler wittenburg2016dynamics , Lagrangian goldstein2002classical , or Hamiltonian mechanics marsden1999introduction . While these classical methods provide accurate and interpretable models, they are challenged in scenarios where the underlying dynamics are partially unknown, highly complex, or subject to unmodeled effects such as friction, contact forces, or material nonlinearities glocker2001set .
@@
Machine learning has opened the door to data-driven simulation of dynamical systems, offering the potential to learn complex behaviors directly from observational data baker2022pde ; karniadakis2021physics . Early approaches in this domain employed classical machine learning architectures such as Long Short-Term Memory (LSTM) networks vlachas2018data ; hochreiter1997long and Fully Connected Neural Networks (FCNNs) sanchez2020learning to learn discrete-time mappings from current states to future states. While these methods have shown promise in short-term prediction tasks, they show limited out-of-distribution generalization. Specifically, these black-box predictors learn fixed-timestep transitions $f:s_{t}\mapsto s_{t+\Delta t}$, which not only limits their temporal resolution but also causes error accumulation that grows exponentially with the prediction horizon ren2025learning . Furthermore, such discrete-time approaches often fail to capture the continuous nature of physical dynamics or enforce fundamental conservation laws of conservative systems zhong2023sympnets .
@@
Neural Ordinary Differential Equations (NODEs) chen2018neural marked a major step in continuous-time learning by representing system dynamics as neural vector fields integrated with standard ODE solvers. This formulation handles irregular time sampling and directly connects learning with established numerical integration theory morrill2021neuralode . This approach has been successfully applied to various physical systems, with extensions including Hamiltonian Neural Networks greydanus2019hamiltonian , Lagrangian Neural Networks cranmer2020lagrangian , and constrained formulations for systems with known invariants xiao2022learning . Despite their strengths, NODEs suffer from a bottleneck: the adjoint sensitivity method for backpropagation requires solving an additional backward ODE, which increases computational cost and memory use, and limits scalability for stiff or high-dimensional systems gholami2019anode ; pal2022chemnode .
@@
Recent advances in flow-matching algorithms liu2022flow ; lipman2023flow show that vector fields can be learned by directly supervising derivatives at sampled points, avoiding backward ODE solves tong2023improving . Applying this idea to dynamical systems, we train neural networks to approximate accelerations directly, eliminating costly integration in the learning loop li2024fourier . We propose Flow-Matching Neural Ordinary Differential Equations (FNODE), a framework for efficient data-driven modeling of multibody systems. Instead of predicting future states via time integration, FNODE trains a neural network to learn directly the acceleration vector field – the natural quantity in mechanical systems governed by force laws arnold1989mathematical . This reformulation removes the need for ODE solvers in backpropagation, improving computational efficiency while preserving accuracy. By learning accelerations directly, the model captures the instantaneous dynamics of the system, leading to better generalization and more stable long-term predictions. Moreover, FNODE naturally accommodates various numerical differentiation schemes for computing acceleration targets from position data, including finite differences and spectral methods, allowing practitioners to balance between noise robustness and accuracy based on their data characteristics fornberg1988generation .
@@
    We evaluate FNODE on a diverse set of multibody dynamics benchmarks, showing consistent improvements over baselines—including Neural ODEs, Hamiltonian Neural Networks, LSTMs, and FCNNs—in long-term prediction accuracy, out-of-distribution generalization, and computational efficiency.
@@ 2.1 The Multibody Dynamics Problem
We employ a MBD formulation in redundant generalized coordinates wherein Lagrange multipliers are used to incorporate constraint equations directly into the equations of motion shabana2013dynamics .
@@ 2.3.2 Finite Difference Method
For non-periodic datasets, finite difference (FD) schemes provide a simple and robust alternative to FFT-based differentiation. Using Taylor expansions, the first derivative can be approximated by Atkinson1989 :
@@ 2.4 Error Propagation in Integration
While FNODE avoids numerical integration during training, its long-term accuracy depends on how errors propagate through the ODE solver. To carry out a basic error analysis following a classical framework HaNW93 ; BrCP96 , let the true system state at step $n$ be $Z_{n}$, advanced by the exact evolution operator $\Phi_{g}$, and let $\hat{Z}_{n}$ denote the predicted state advanced by the learned solver $\Phi_{f}$. The global error is $E_{n}=\hat{Z}_{n}-Z_{n}$. The exact and predicted updates are
@@ 3.3 Triple-Mass-Spring-Damper System
$\mathbf{m_{1}}$$\mathbf{k_{1}}$$\mathbf{x}$$\mathbf{d_{1}}$$\mathbf{m_{2}}$$\mathbf{k_{2}}$$\mathbf{d_{2}}$$\mathbf{m_{3}}$$\mathbf{k_{3}}$$\mathbf{d_{3}}$
@@ 3.4 Double Pendulum
$\mathbf{x}$$\mathbf{y}$$\mathbf{l_{1}}$$\mathbf{m_{1}}$$\mathbf{l_{2}}$$\mathbf{m_{2}}$$\mathbf{\theta_{1}}$$\mathbf{\theta_{2}}$
@@ Appendix A Algorithms for training the FNODE
1:Initialize: FNODE network $f(\cdot,\bm{\Theta})$ predicting accelerations
2:Input: Ground truth state-acceleration pairs $\mathcal{D}=\{(\mathbf{S}_{i},\mathbf{a}_{i})\}_{i=0}^{T-1}$ where $\mathbf{S}_{i}\in\mathbb{R}^{2n_{b}}$ and $\mathbf{a}_{i}\in\mathbb{R}^{n_{b}}$, optimizer and its settings
@@
4:  for each training sample $(\mathbf{S}_{i},\mathbf{a}_{i})\in\mathcal{D}$ do
5:   Prepare input state $\mathbf{S}_{i}=[\mathbf{s}_{i}^{\text{pos}},\mathbf{s}_{i}^{\text{vel}}]$ and target acceleration $\mathbf{a}_{i}$
6:   Predict Acceleration: $\hat{\mathbf{a}}_{i}=f(\mathbf{S}_{i},\bm{\Theta})$
7:   Compute loss $L=\|\mathbf{a}_{i}-\hat{\mathbf{a}}_{i}\|_{2}^{2}$
@@
11:Output: Trained FNODE $f(\cdot,\bm{\Theta}^{*})$
@@
1:Initialize: FNODE network $f(\cdot,\bm{\Theta})$ for minimal coordinates predicting accelerations; Choose integrator $\Phi$, uses prior knowledge of constraint equation $\phi$ and the minimal coordinates.
2:Input: Minimal coordinate state-acceleration pairs $\mathcal{D}^{M}=\{(\mathbf{S}_{i}^{M},\mathbf{a}_{i}^{M})\}_{i=0}^{T-1}$ where $\mathbf{S}_{i}^{M}\in\mathbb{R}^{2m}$ and $\mathbf{a}_{i}^{M}\in\mathbb{R}^{m}$, constraint equations $\varphi$, optimizer and its settings
@@
4:  for each training sample $(\mathbf{S}_{i}^{M},\mathbf{a}_{i}^{M})\in\mathcal{D}^{M}$ do
5:   Prepare minimal state $\mathbf{S}_{i}^{M}=[\mathbf{s}_{i}^{M,\text{pos}},\mathbf{s}_{i}^{M,\text{vel}}]$ and target acceleration $\mathbf{a}_{i}^{M}$
6:   Predict Minimal Acceleration: $\hat{\mathbf{a}}_{i}^{M}=f(\mathbf{S}_{i}^{M},\bm{\Theta})$
7:   Compute minimal loss $L=\|\mathbf{a}_{i}^{M}-\hat{\mathbf{a}}_{i}^{M}\|_{2}^{2}$
@@
11:Output: Trained FNODE $f(\cdot,\bm{\Theta}^{*})$ for minimal coordinates

### v2

@@ 1 Introduction
Multibody dynamics simulation is relevant in engineering design and scientific computation, with applications in numerous fields, e.g., robotics featherstone2008rigid , vehicle and train dynamics shabana2020dynamics , biomechanics delp2007opensim , and molecular dynamics schlick2010molecular , to name a few. Traditionally, modeling multibody systems has relied on first-principles approaches such as Newton-Euler wittenburg2008dynamics , Lagrangian goldstein2002safko , or Hamiltonian mechanics marsden1999introduction . While these classical methods provide accurate and interpretable models, they are challenged in scenarios where the underlying dynamics are partially unknown, highly complex, or subject to unmodeled effects such as friction, contact forces, or material nonlinearities glocker2013set .
@@
Machine learning has opened the door to data-driven simulation of dynamical systems, offering the potential to learn complex behaviors directly from observational data raissi2019physics ; karniadakis2021physics . Early approaches in this domain employed classical machine learning architectures such as Long Short-Term Memory (LSTM) networks vlachas2018data ; hochreiter1997long and Fully Connected Neural Networks (FCNNs) sanchez2020learning to learn discrete-time mappings from current states to future states. While these methods have shown promise in short-term prediction tasks, they show limited out-of-distribution generalization. Specifically, these black-box predictors learn fixed-timestep transitions $f:s_{t}\mapsto s_{t+\Delta t}$, which not only limits their temporal resolution but also causes error accumulation that grows exponentially with the prediction horizon ren2024learning . Furthermore, such discrete-time approaches often fail to capture the continuous nature of physical dynamics or enforce fundamental conservation laws of conservative systemsjin2020sympnets .
@@
Neural Ordinary Differential Equations (NODEs) chen2018neural marked a major step in continuous-time learning by representing system dynamics as neural vector fields integrated with standard ODE solvers. This formulation handles irregular time sampling and directly connects learning with established numerical integration theory kidger2020neural . This approach has been successfully applied to various physical systems, with extensions including Hamiltonian Neural Networks greydanus2019hamiltonian , Lagrangian Neural Networks cranmer2020discovering , and constrained formulations for systems with known invariantsyang2020learning . Despite their strengths, NODEs suffer from a bottleneck: the adjoint sensitivity method for backpropagation requires solving an additional backward ODE, which increases computational cost and memory use, and limits scalability for stiff or high-dimensional systems gholaminejad2019anode ; wang2024mbdnode .
@@
Recent advances in flow-matching algorithms liu2023flow ; lipman2023flow show that vector fields can be learned by directly supervising derivatives at sampled points, avoiding backward ODE solves tong2024improving . Applying this idea to dynamical systems, we train neural networks to approximate accelerations directly, eliminating costly integration in the learning loop li2020fourier . We propose Flow-Matching Neural Ordinary Differential Equations (FNODE), a framework for efficient data-driven modeling of multibody systems. Instead of predicting future states via time integration, FNODE trains a neural network to learn directly the acceleration vector field – the natural quantity in mechanical systems governed by force laws arnol1997mathematical . This reformulation removes the need for ODE solvers in backpropagation, improving computational efficiency while preserving accuracy. By learning accelerations directly, the model captures the instantaneous dynamics of the system, leading to better generalization and more stable long-term predictions. Moreover, FNODE naturally accommodates various numerical differentiation schemes for computing acceleration targets from position data, including finite differences and spectral methods, allowing practitioners to balance between noise robustness and accuracy based on their data characteristics fornberg1988generation .
@@
    We evaluate FNODE on a diverse set of multibody dynamics benchmarks, showing consistent improvements over baselines—including MBDNODEs, LSTMs, and FCNNs—in long-term prediction accuracy, out-of-distribution generalization, and computational efficiency.
@@ 2.1 The Multibody Dynamics Problem
We employ a MBD formulation in redundant generalized coordinates wherein Lagrange multipliers are used to incorporate constraint equations directly into the equations of motion shabana2020dynamics .
@@ 2.3.2 Finite Difference Method
For non-periodic datasets, finite difference (FD) schemes provide a simple and robust alternative to FFT-based differentiation. Using Taylor expansions, the first derivative can be approximated by atkinson2008introduction :
@@ 2.4 Error Propagation in Integration
While FNODE avoids numerical integration during training, its long-term accuracy depends on how errors propagate through the ODE solver. To carry out a basic error analysis following a classical framework hairer1993classical ; brenan1996numerical , let the true system state at step $n$ be $Z_{n}$, advanced by the exact evolution operator $\Phi_{g}$, and let $\hat{Z}_{n}$ denote the predicted state advanced by the learned solver $\Phi_{f}$. The global error is $E_{n}=\hat{Z}_{n}-Z_{n}$. The exact and predicted updates are
@@ Appendix A Algorithms for training the FNODE
1:Initialize: FNODE network $f(\cdot;\bm{\Theta})$ predicting accelerations
2:Input: Ground truth state–acceleration pairs $\mathcal{D}=\{(\mathbf{Z}_{i},\ddot{\mathbf{z}}_{i})\}_{i=0}^{T-1}$ where $\mathbf{Z}_{i}\in\mathbb{R}^{2n_{b}}$ and $\ddot{\mathbf{z}}_{i}\in\mathbb{R}^{n_{b}}$, optimizer and its settings
@@
4:  for each training sample $(\mathbf{Z}_{i},\ddot{\mathbf{z}}_{i})\in\mathcal{D}$ do
5:   Prepare input state $\mathbf{Z}_{i}=[\mathbf{z}_{i},\dot{\mathbf{z}}_{i}]$ and target acceleration $\ddot{\mathbf{z}}_{i}$
6:   Predict Acceleration: $\hat{\ddot{\mathbf{z}}}_{i}=f(\mathbf{Z}_{i};\bm{\Theta})$
7:   Compute loss $L=\|\ddot{\mathbf{z}}_{i}-\hat{\ddot{\mathbf{z}}}_{i}\|_{2}^{2}$
@@
11:Output: Trained FNODE $f(\cdot;\bm{\Theta}^{*})$
@@
1:Initialize: FNODE network $f(\cdot;\bm{\Theta})$ for minimal coordinates predicting accelerations; Choose integrator $\Phi$, use prior knowledge of constraint equation $\varphi$ and the minimal coordinates.
2:Input: Minimal coordinate state–acceleration pairs $\mathcal{D}^{M}=\{(\mathbf{Z}_{i}^{M},\ddot{\mathbf{z}}_{i}^{M})\}_{i=0}^{T-1}$ where $\mathbf{Z}_{i}^{M}\in\mathbb{R}^{2m}$ and $\ddot{\mathbf{z}}_{i}^{M}\in\mathbb{R}^{m}$, constraint equations $\varphi$, optimizer and its settings
@@
4:  for each training sample $(\mathbf{Z}_{i}^{M},\ddot{\mathbf{z}}_{i}^{M})\in\mathcal{D}^{M}$ do
5:   Prepare minimal state $\mathbf{Z}_{i}^{M}=[\mathbf{z}_{i}^{M},\dot{\mathbf{z}}_{i}^{M}]$ and target acceleration $\ddot{\mathbf{z}}_{i}^{M}$
6:   Predict Minimal Acceleration: $\hat{\ddot{\mathbf{z}}}_{i}^{M}=f(\mathbf{Z}_{i}^{M};\bm{\Theta})$
7:   Compute minimal loss $L=\|\ddot{\mathbf{z}}_{i}^{M}-\hat{\ddot{\mathbf{z}}}_{i}^{M}\|_{2}^{2}$
@@
11:Output: Trained FNODE $f(\cdot;\bm{\Theta}^{*})$ for minimal coordinates